I have came to this. The sum of two numbers is 82. The second is 6 more than 3 times the first. What are the two numbers?
I got loat when I got to this point can anyone help me

Answer by wuwei96815(245)

Let x=the first number
Then 3x+6=the second number

x + (3x+6) = 82
x + 3x +6 = 82
4x = 82 - 6 = 76
4x = 76
x = 76/4 = 19

3x+6 = 63

We know we are right because 19 and 63 = 82. Doesn't it?